OpenStudy (anonymous):

What is the solution to the equation log3^x – log3^2 = 2

OpenStudy (anonymous):

Is it log base 10?

OpenStudy (zepp):

\(\large \log3^x-log3^2=2\) Logarithm super properties \(\large \log3^{x\div 2}=2\)

OpenStudy (anonymous):

im not sure its either 3 or 10. the question isnt clear

OpenStudy (zepp):

Write it as exponential form \(\large 3^2 = x \div 2\) Can you take it from there?

OpenStudy (anonymous):

answer are x=9/4 x=4/3 x=18 and x=12

OpenStudy (anonymous):

don't think so

OpenStudy (zepp):

\(3^2=x \div 2\) \(9 = x\div 2\) \(x = 2*9\)

OpenStudy (zepp):

Owait, in fact

OpenStudy (zepp):

\((x \div 2)log3 =2\)

OpenStudy (zepp):

Should be like this

OpenStudy (anonymous):

thats x/2 right?

OpenStudy (zepp):

Then \(\large \log3 = \frac{2}{(x \div 2)}\)

OpenStudy (dumbcow):

um is this log base 3 we are talking about ? \[\log_{3} x - \log_{3}2 = 2 \] OR \[\log (3^{x})-\log(3^{2}) = 2\]

OpenStudy (dumbcow):

based on answer choices, im assuming base 3 then zepp is correct , x = 18

OpenStudy (zepp):

\(\large \log3 = \frac{4}{x}\) I can't get a rational answer with exponent :(

OpenStudy (anonymous):

its cool its only worth 3 points but thanks for the help. appreciate it

OpenStudy (dumbcow):

hello? what base is the log

OpenStudy (anonymous):

either 10 or 3. the teacher sorta sucks, why im asking here instead of doing it on my own

OpenStudy (dumbcow):

it has nothing to do with the teacher...on your paper or book how is it written, is the 3 a subscript or not ?

OpenStudy (anonymous):

agreed, the teacher now days are overpaid and underworked

OpenStudy (anonymous):

log 3 x – log 3 2 = 2 thats how it looks

OpenStudy (dumbcow):

teachers overpaid ?? are you crazy

OpenStudy (dumbcow):

not in USA

OpenStudy (dumbcow):

anyway looks like base 3 --> x = 18

OpenStudy (anonymous):

yup thats what zepp said

OpenStudy (zepp):

I'm awesome. :D